What is the age distribution of patients who make office visits to a doctor or nurse?

The following table is based on information taken from a medical journal.

Age group, years Under 15 15-24 25-44 45-64 65 and older

Percent of office visitors 15% 10% 20% 15% 40%

Suppose you are a district manager of a health management organization (HMO) that is monitoring the office of a local doctor or nurse in general family practice. This morning the office you are monitoring has eight office visits on the schedule. What is the probability of the following?

(b) From 2 to 5 patients are 65 years old or older (include 2 and 5). (Round your answer to three decimal places.)

(c) From 2 to 5 patients are 45 years old or older (include 2 and 5). (Hint: Success if 45 or older. Use the table to compute the probability of success on a single trial. Round your answer to three decimal places.)

(e) All the patients are 15 years old or older. (Round your answer to three decimal places.)
Answer #1

b)

here this is binomial with parameter n=8 and p=0.4
P(2<=X<=5)= x=ab     (nCx)px(1−p)(n-x)    = 0.844

c)

this is binomial with parameter n=8 and p=0.55
P(2<=X<=5)= x=ab     (nCx)px(1−p)(n-x)    = 0.762

d)

this is binomial with parameter n=8 and p=0.85
P(X=8)= (nCx)px(1−p)(n-x)    = 0.272